OpenStudy (anonymous):

Mary applies for a loan with a 7/9 percent interest rate. what is 7/9 percent as a decimal?

OpenStudy (anonymous):

If you want to convert percent to decimal, divide by 100. \[\frac{7}{9}\% = \frac{7/9}{100}= \frac{7}{900} = ~?\] I'm sure you can do the last bit to get the answer